Meeting Minutes 1st August 2013
Minutes from 1/08
Attendees :- Stevie Conlon, Neil Grant, Michael Sun, Richard Good , Dave Berry
Apologies : Tim Gray, Peter Jackson.
Next meeting 08/08 4.09 at 3:00.
Update on Current Tasks
- TEST server work nearing completion, Fault handling and Test suite to be deployed. New Email created for deployment use by service management. Authentication issue with WS.
- LIVE server work in progress. Fault handling and Test suite to be deployed. New Email created for deployment use by service management.
- DR request of ITI, see action 22. DR test did not take place on the 18th as planned to be rescheduled.
- Additional days required for standby database creation as part of new infrastructure for TEST and LIVE.- DBs created .
Actions
16. Outstanding JIRA to be closed and completed by Dev Tech Team. Preferred for build sign off.
Action : Richard/Michael – complete for Development Team.
Action Neil : 2 outstanding JIRAs with Dev Tech
Update on 26th. Ongoing to be reviewed as part of the build review.
Update on the 10th. Solutions to outstanding JIRAs to be implemented and tested.
Update on the 18th. Two outstanding JIRA.
17. Arrange build review and sign off
Action: Stevie
Update for 12th May. Awaiting TAD. To arrange w/c 17th.
Update for the 26th . Build Review arranged for the 3rd July. PPIR by email for 4th July.
Update for the 3rd. PBBR not signed and TAD to be completed and JIRA solutions implemented.
New milestones to be planned.
22. Peter to initiate DR test with ITI involvement. As the service is not resilient, this will involve bringing down the current test server and bringing it up elsewhere. This will require around 1 days downtime for the test server..
Action Peter/DevTech
Update for the 10th. DR test schedule with ITI for the 18th July on the SOA TEST server.
Update on the 18th. Pulled on the day. To be rescheduled with ITI.
Update on the 25th . Stevie/Dev Tech to re-arrange with ITI
Update on the 1st. No response yet from ITI.
26. Peter to test SOA Dehydration Database Housekeeping scripts on Friday 28th.
Action Peter
Update for 10th July. Peter prepped for testing on Friday 28th. Peter has been testing the scripts on the Dev database.
Update for 18th. Peter has tested the scripts on Dev and passed to Neil for handover to Production Management. Complete.
Update for 25th. Stevie to follow up with Peter to move JIRA and docs to Neil.
Update fofr the 1s . JIRA moved to Neil to complete
28. Deploy test suite and fault handling suite successfully to TEST environments.
Action Neil/Michael
Update for the 1st August. Error persists with the authorisation for java web services.
Closed Actions
21. Determine Tech Man Involvement. Arrange discussion with Dev Tech and Tech Man to allow Heather L to determine involvement as no running applications will be on the servers at present.
Action Stevie, Dev Tech,Tech Man.
Update required for 26th. See action 24. Closed
23. Report potential issue to Oracle (as described above)
Action Michael.
Update required for 25th June. Complete .Call raised with Oracle
24,Meet with Tech Man ( Heather Larnach) to verify their requirements from the project.
Action : Stevie
Update. Stevie met with Heather who will allocate resource to review the TAD and servers with Dev Tech. Closed
8. TAD to be updated ongoing with completion expected w/c 06/05 . To be Signed Off at Dev tech weekly Meeting , 09/05. TAD to contain policy of SOA DB and log archiving limit of 60 days. Required for build sign off.
Action Neil :
Update on 5th :- TAD to complete when all known tasks resolved on TEST.
Update on the 26th . TAD to be circulated Thursday 27th for review.
Update on the 11th TAD to be reviewed at the Dev Tech meeting on the 12th July.
26. Peter to test SOA Dehydration Database Housekeeping scripts on Friday 28th.
Action Peter
Update for 10th July. Peter prepped for testing on Friday 28th. Peter has been testing the scripts on the Dev database.
Update for 17th. Peter has tested the scripts on Dev and passed to Neil for handover to Production Management. Complete.
25. Follow up outstanding Oracle call with Oracle contact
Action Stevie/Michael
Update for 10th July. Query has been passed to Oracle contact. The current solution is acceptable but Oracle may be able to suggest a better way in the future.
Update for 18th, still with Oracle.
Update for the 25th. If Oracle come back with a different way of completing . this will be reviewed. JIRA closed ,. Action Closed
27. Neil to meet with Tech Management to determine DR testing.
Action Neil/Tech Man.
Update for 1st August. Meeting held and outcome documented. Action Closed